convert decimals to % thanks
0.40=
1.20=
0.05=
0.001/2=
0.375=
percents to fractions 60%=
7%=

Answer by bucky(2097)

Convert to percent
.
Do this by multiplying by 100 and adding a percent sign. Note that multiplying by 100 moves
the decimal point two places to the right:
.
0.40 ... 4.*100 = 40%
.
1.20= ... 1.20*100 = 120%
.
0.05= ... .05*100 = 5%
.
0.001/2= 0.0005 .... 0.0005*100 = 0.05%
.
0.375= .... 0.375*100 = 37.5%
.
The next set of conversions is in the reverse direction, converting percents to fractions
.
Do this by dropping the percent sign and place the number over 100
.
60%= 60/100 and this simplifies (by dividing numerator and denominator by 20) to 3/5
.
7%= 7/100
.
Hope this helps you to do this type of conversion ... decimals to percent and from percent
to fractions.